A car generating a power of 97680.67watts is traveling on a road with a coefficient of kinetic friction of 0.051. If the velocity of it upon reaching its destination is 30.33m/s and has been accelerating at a rate of 3.56m/s2, what is the mass of the car?
1
Expert's answer
2020-11-25T10:46:33-0500

Power can be given as

"P=Fv"

And "F=ma"

"P=mav\\m=\frac{P}{av}"


"m=\frac{97680.67}{30.33\times 3.56}\\m=904.6kg"


So mass of car is 904.6 kg